Sol Summary Report – February 20th

Crew 275 Sol Summary Report 20Feb2023

Sol: 8

Summary Title: Sol 8 – Exploring the eighth continent

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today we conducted the first EVA dedicated to photogrammetry. It started very successfully with a stop at the location of the atmospheric instruments to retrieve some data, change batteries, and conduct supplementary tests on Mega-Ares. The EVA crew then reached North Ridge to start their activities: sample collection and 3D mapping of the terrain using drones. If the scientific objectives were completely fulfilled, it is important to make a few remarks regarding this EVA. 1) One crewmember had a problem with their hat and headset during the operations. There was a misunderstanding regarding the actions to take in this situation, which led to a few minutes before taking the decision to put the hand inside the helmet to solve the issue. This event was discussed as a group and the procedure is now clear for everyone. 2) There were issues regarding radio communications within the crew and between HabCom and the Crew. This led to some confusion. Again, this event was debriefed and the procedure is also clear for everyone now. It includes some reminders on how to communicate efficiently on the different radio channels. 3) The EVA crew was approached by 4 pick-ups getting as close as 150 m from them. They took pictures of the crew conducting their activities. This was reported as a very stressful event by the EVA crew, as this took a lot of their mental space as they were piloting a drone at this moment.

Look Ahead Plan:

The tests conducted during the last two EVAs have shown that there was a small issue with the antenna of Mega-Ares, which can be solved only from within the Hab. For this reason, we would like to perform an EVA tomorrow (Sol 9) morning to retrieve this antenna and proceed to a few additional checks on the atmospheric instruments. We plan to put this antenna back in its place on Wednesday (Sol 10) afternoon and to take this opportunity to change the batteries of all the atmospheric instruments.

With the success of the first photogrammetry EVA, a 3D map of the target area has been rendered. Now, we would like to send two EVA crews to the exact same location to check the added value of the 3D map to ease the ability of astronauts to identify some specific features onsite. We thus plan to send an EVA crew to North Ridge with a 3D map on Wednesday morning (Sol 10) and another one with only a 2D map on Thursday morning (Sol 11).

Anomalies in work: None.

Weather: Sunny, low -8°C/18°F, high 6°C/42°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site